### v1

# Intrinsic Dimensionality as a Model-Free Measure of Class Imbalance

Abstract
Imbalance in classification tasks is commonly quantified by the cardinalities of examples across classes. This, however, disregards the presence of redundant examples and inherent differences in the learning difficulties of classes. Alternatively, one can use complex measures such as training loss and uncertainty, which, however, depend on training a machine learning model. Our paper proposes using data Intrinsic Dimensionality (ID) as an easy-to-compute, model-free measure of imbalance that can be seamlessly incorporated into various imbalance mitigation methods. Our results across five different datasets with a diverse range of imbalance ratios show that ID consistently outperforms cardinality-based re-weighting and re-sampling techniques used in the literature. Moreover, we show that combining ID with cardinality can further improve performance. Code: https://github.com/cagries/IDIM.
keywords:
intrinsic dimension , long-tailed visual recognition , class imbalance , long-tailed learning
journal: Preprint submitted to arXiv
[…]
[ceng]organization=Department of Computer Engineering, addressline=Middle East Technical University, city=Ankara, postcode=06810, state=Ankara, country=Turkey
[…]
1 Introduction
Learning-based visual recognition is prone to differences among classes in training datasets. One aspect of such a difference is the cardinalities of examples across classes (called class cardinalities in short in the paper): Datasets generally exhibit a long-tailed distribution in terms of class cardinalities [1, 2, 3, 4], which has prompted research using class cardinality as a measure in various imbalance mitigation strategies [5, 6, 7] – see also Fig. 1a.
Figure 1: Existing approaches to long-tailed visual recognition rely on (a) the cardinalities of examples across classes, which is affected by redundant examples in the dataset, or (b) measures of class hardness, which require training an ML model. (c) Our approach of using data ID provides information about the manifold (and hence, complexity) of each class without being affected by redundant examples and without training an ML model. (d) We show that this simple, model-free, plug-and-play approach provides significant improvements with different approaches (a sample result shown here with a resampling approach).
Another aspect of imbalance that is more relevant to the learning dynamics involves the inherent discriminability or learning hardness of classes [8, 4, 9, 10]. Although such hardness-based measures can be viewed as more established and promising, they require multiple training stages since a machine learning model needs to be trained first to quantify imbalance among classes (Fig. 1b).
Figure 2: Estimated IDs for CIFAR10 classes for different sample counts. ID can capture inherent differences among classes and is relatively robust against sample count, as exemplified here with the CIFAR10 dataset, subsampled at different sample counts per class.
In this paper, we introduce a model-free, cardinality-free measure of imbalance among classes based on data Intrinsic Dimensionality (ID) (Fig. 1c). ID estimation [11, 12, 13] is a well-established field of research that aims to estimate the implicit dimensions of lower-dimensional data manifolds embedded in a higher-dimensional space.
We first show that class-wise calculation of ID can point to inherent differences among classes (see, e.g., Fig. 2). For a given dataset, it is sufficient to compute the ID for each class once. Then, we integrate our ID-based measure into various imbalance mitigation strategies based on resampling, loss reweighting, and margin adjustment methods. We then compare ID-based imbalance mitigation against alternative approaches and evaluate ID-enhanced training combined with state-of-the-art methods.
Why Data ID? Manifold Learning Theory suggests that high-dimensional data often lie on a lower-dimensional manifold [14, 15, 16, 17]. Our results confirm that, in a multi-class dataset, the data manifold of each class has a specific ID [18]. Data ID quantifies this true dimensionality. Statistical learning theory [19] suggests that the number of effective samples needed for a target generalization error scales with (true) data dimensionality. Thus, Data ID for a class – being independent of its sample size – can act as a proxy for the amount of data required for modeling that class [12]. Instead of relying only on the cardinality of classes, which completely disregards intrinsic properties of data, we argue that IDs of classes better capture the underlying structure of the data, leading to increased robustness in imbalance mitigation strategies, as we experimentally demonstrate. Furthermore, our analysis on ID estimation (Section 4) shows that data ID is robust against sample count, sample noise and extrinsic dimension, qualities that we would expect from a measure of imbalance for long-tailed data distributions.
Main Contributions. (1) We introduce a novel measure of class imbalance based on data ID. We view this as a new perspective into the class imbalance literature as it complements existing measures based on cardinality and class hardness. (2) We show that our ID-based measure can be easily integrated into different imbalance mitigation strategies. (3) We demonstrate that our ID-based measure is able to quantify semantic imbalance among classes, which is missed by cardinality-based methods. (4) Across different datasets, we report significant improvements over using class cardinality as a measure. For example, incorporating data ID into progressively-balanced sampling increases accuracy by up to $2.1\%$ (Fig. 1(d)) with virtually zero additional cost.
2 Related Work
2.1 Long-tailed Visual Recognition
Long-tailed recognition approaches mitigate the effects of imbalance relying on sample counts or hardness under the following main categories [2, 1, 20]: Resampling methods [7, 6], loss-reweighting methods [8, 9, 21, 22], margin (logit) adjustment methods [23, 24] and other methods [25, 26, 27]. For a detailed review, see A.1 and, for formal definitions of the common approaches, see Section 5.
2.2 Intrinsic Dimensionality
Intrinsic dimensionality analysis focuses on quantifying the intrinsic qualities of data. As a general definition, the intrinsic dimension is the minimum number of parameters required to represent a data distribution with minimal loss of generality [11]. ID is closely related to the manifold hypothesis [14], which posits that high dimensional natural image datasets lie on a lower dimensional manifold, enabling learning. In this context, ID refers to the number of dimensions of this lower-dimensional manifold. The ID of a dataset can be estimated globally (i.e., for the entire dataset), or locally (around a local neighborhood), depending on the scope and the estimation method. ID estimation methods in the literature can be studied in two general categories: Projective methods [28, 29, 30, 31] and geometric (topological) methods [32, 33, 34]. For the sake of space, we provide a further review of these methods in A.2.
A recent work by Ma et al. [18] focuses on the intrinsic dimension of perceptual manifolds of balanced datasets in the context of model fairness and introduces ID regularization as an end-to-end trainable loss, and similarly the work by Ansuini et al. [35] also studies the ID of the representation of data in the intermediate layers of a neural network. In comparison, we introduce ID as an imbalance measure and show that it can be used for training as a model-free mitigation method in the context of imbalanced datasets.
2.3 Comparative Summary
We observe that the class imbalance literature has mainly used class cardinality or class hardness (obtained through some measure of learning difficulty, e.g., prediction probability, class loss). To our knowledge, we are the first to introduce data ID as a model-free measure of class imbalance and use it in various imbalance mitigation strategies.
3 FisherS: Background on ID Estimation
We first introduce the (data) Intrinsic Dimension (ID) estimation method we have used, the method of Albergante et al. [29], which we call FisherS, for estimating the IDs of individual classes. As we will analyze through experimentation, our approach and conclusions are not dependent on the specific ID estimation method.
In essence, FisherS utilizes the concept of blessing of dimensionality where points from samples with high dimensionality can be separated from the rest of the sample set by linear inequalities with high probability, regardless of the sample count [14]. With some preprocessing, these linear inequalities can be estimated by Fisher’s linear discriminant in an efficient manner.
For a dataset $X$ with points $\mathbf{x}\in\mathbb{R}^{n}$, the FisherS ID is estimated with the following steps:
Preprocessing. The data is initially preprocessed by centering and projecting onto the linear subspace spanned by its first major principal components found by PCA [36] to reduce collinearity in the data. The major components are defined as components whose eigenvalues are higher than $\frac{\lambda_{max}}{C}$, with $\lambda_{max}$ being the largest eigenvalue of the sample covariance matrix, and $C$ a constant (a good value being $C=10$ to reduce collinearity [14, 37]). Afterwards, whitening (i.e., linearly transforming data points such that the covariance matrix is an identity matrix) is applied, and the obtained vectors are normalized to unit length so that each vector is a projection into a unit sphere.
Fisher Separability. Given a dataset preprocessed as described above, a point $\mathbf{x}\in X$ is said to be Fisher-separable from a collection of points $Y$ with parameter $\alpha\in[0,1)$ if:
$\langle\textbf{x},\textbf{y}\rangle\leq\alpha\langle\textbf{x},\textbf{x}\rangle,$ (1)
for all $\mathbf{y}\in Y$, with $\langle\mathbf{x},\mathbf{y}\rangle=\sum^{n}_{k=1}x_{k}y_{k}$ being the inner product of $\mathbf{x},\mathbf{y}\in\mathbb{R}^{n}$ [14]. Similarly, the dataset $X$ is Fisher-separable with parameter $\alpha$ if (Eq. 1) is true for all $\mathbf{x}\in X$.
FisherS ID Estimation. Let $p_{\alpha}(\mathbf{y})$ be the probability that a point $\mathbf{y}$ is not separable from all other points, and $\bar{p_{\alpha}}(\mathbf{y})$ be the mean value of the distribution of $p_{\alpha}(\mathbf{y})$ over all data points. With the data projected onto a unit sphere, Albergante et al. [29] show that $p_{\alpha}(\mathbf{y})$ can be estimated on the rotationally invariant equidistribution on the unit sphere $\mathbb{S}^{n-1}\subset\mathbb{R}^{n}$ by:
$p_{\alpha}\approx\frac{1}{\alpha\sqrt{2\pi n}}(1-\alpha^{2})^{\frac{n-1}{2}},$ (2)
where the distribution of $p_{\alpha}$ is a delta function centered in $\bar{p_{\alpha}}$. In this case, the ID $n_{\alpha}$ of the data for parametrized $n_{\alpha}$ can be obtained by solving for $n$ in (2) as:
$n_{\alpha}=\frac{1}{-\ln(1-\alpha^{2})}W\left(\frac{-\ln(1-\alpha^{2})}{2\pi\bar{p}_{\alpha}^{2}\alpha^{2}(1-\alpha^{2})}\right),$ (3)
where $W(\cdot)$ is the Lambert function. In FisherS, this process is repeated for many values of $\alpha$, and the $\alpha$ parameters with valid $n_{\alpha}$ dimensions are kept in $\boldsymbol{\alpha}_{s}$. The $n_{\alpha}$ corresponding to $\alpha$ that is closest to $0.9\times\max\{\boldsymbol{\alpha}_{s}\}$ is determined to be the intrinsic dimension, $d$, of the dataset $X$.
Our Application. In our study, we estimate class-wise data ID for each class $c$, denoted as $d_{c}$, by applying the aforementioned FisherS ID estimation to the (training) samples of class $c$. Note that we do this once per dataset, without requiring any form of deep model training beforehand. Our method relies on the relative differences of ID estimates between classes. As a result, for our mitigation experiments, we normalize $d_{c}$ to obtain $\hat{d}_{c}$ so that $\sum_{c}\hat{d}_{c}=1$:
$\hat{d}_{c}=\frac{\textrm{FisherS}(\{\mathbf{x}_{i}\ |\ y_{i}=c\})}{\sum_{c^{\prime}}\textrm{FisherS}(\{\mathbf{x}_{i}\ |\ y_{i}=c^{\prime}\})}.$ (4)
(a) ID vs. sample count [synthetic data].
(b) ID vs. extrinsic dimension [synthetic data].
Figure 3: Analysis on FisherS estimated IDs. The analysis with the synthetic data in (a) and a real dataset in Fig. 2 suggest that ID estimation is robust against sample count. Moreover, the analysis in (b) shows that it is not affected by extrinsic dimensionality either. Dashed lines: true ID values.
4 ID as a Measure of Class Imbalance
Why an ID-based imbalance measure? Class imbalance analysis and mitigation methods require quantifying the amount of imbalance across classes. As discussed in Introduction, data ID by definition provides a measure of the true dimensionality (complexity) of the data manifold. Statistical learning theory [19] suggests that the number of effective samples needed for a target generalization error scales with (true) data dimensionality (complexity). Thus, data ID for a class – being independent of its sample count – captures the complexity (true dimensionality) of that class and as suggested by statistical learning theory, provides a proxy for the necessary number of samples required for modeling (learning) that class [12]. This motivates and justifies our use of Data ID with resampling or reweighting methods. Similarly, a higher ID implies a more complex manifold, potentially requiring a decision boundary with wider margins, which justifies using Data ID in logit adjustment. These arguments align with prior work on sample requirements in imbalanced settings, such as the manifold-based derivation of the “effective number of samples” [4] and the feature covariance based measure [17].
ID estimation is robust against sample count. Following similar analysis [38], we generate synthetic data by sampling from a multivariate normal distribution. For ID = $d$, we sample from $\mathcal{N}(\boldsymbol{\mu},\mathbf{\Sigma})$, where $\boldsymbol{\mu}$ is a $d$-dimensional zero vector and $\mathbf{\Sigma}$ is a $d\times d$ identity covariance matrix. Our results in Fig. 3(a) indicate that FisherS ID is robust against changing sample count, which is very important when being used with long-tailed data distributions. Furthermore, in Fig. 2 we calculated the ID estimates of classes on randomly sampled subsets of the real-world CIFAR-10 dataset and observed the resulting ID estimates are relatively robust under a variety of example counts, which reflects this quality on real-world data.
ID estimation is robust against extrinsic dimensionality. For this, we follow a similar approach to Hein and Audibert [38] and generate synthetic data from multivariate normal distributions with fixed intrinsic dimensions, and progressively increase the extrinsic dimension of the data. For ID = $d$, we sample from $\mathcal{N}(\boldsymbol{\mu},\mathbf{\Sigma})$, where $\boldsymbol{\mu}$ is a $D$-dimensional zero vector and $\mathbf{\Sigma}$ is a $D\times D$ covariance matrix with
$\mathbf{\Sigma}=\left[\begin{array}[]{ c c }I_{d\times d}&0\\ 0&0\end{array}\right]_{D\times D}.$ (5)
After sampling, we rotate the generated samples around the origin [39] by applying pairwise rotations for each consecutive pair of dimensions, eliminating constant valued dimensions. Fig. 3(b) shows that ID estimation is robust to changes in extrinsic dimension.
ID estimation is robust against data noise. To investigate the effect of noise for estimating ID in real-world high-dimensional data, we added Gaussian noise with varying scales to CIFAR-10 images (channel intensities normalized between $[0,1]$) and calculated their class-wise ID. Fig. 4 indicates that ID estimation is robust against even extreme levels of noise where the images are hardly identifiable.
Figure 4: ID estimates of CIFAR-10 classes with added Gaussian noise of varying scales ($\sigma\in[0,1]$), showing robustness of ID to even drastic amounts of noise.
ID estimation is robust against correlation. We also show that ID estimation is relatively robust against dimension correlation. See D.4 for this analysis.
5 Using ID to Mitigate Class Imbalance
In this section, we describe how we use per-class ID estimates, $\hat{d}_{c}$ (Eq. 4), in existing imbalance mitigation methods. The underlying principle is that the intrinsic dimension of more complex classes is expected to be higher than that of simpler classes, which should be easier to learn.
Specifically, for resampling-based experiments, we use $\hat{d}_{c}$ for a dataset to specify the sampling probabilities for class $c$, and then uniformly sample from the samples of class $c$. The higher the estimated ID is for a specific class category, the more “difficulty” it poses for a trained model to learn. Therefore we propose that this class should be sampled more frequently for the trained model to capture the intrinsic qualities of the class more effectively. Similarly, when plugging ID-based learning into reweighting methods, the intuitive idea is to have lower loss weighting for “simpler” classes (i.e., classes with lower ID estimates). We do not make use of any sample-size assumptions here and completely rely on the intrinsic qualities captured by ID in this approach.
Resampling methods. In standard sampling, each sample in the training dataset has an equal probability to be sampled, i.e., sampling from a class is directly correlated with the class cardinality of that class [1, 2, 40]:
$p_{\mathbf{x}}=p_{c}\times\frac{1}{N_{c}},$ (6)
where $p_{\mathbf{x}}$ is the probability of sampling a training sample $\mathbf{x}$ of class $c$ and $N_{c}$ the number of samples in class $c$. Resampling methods define sampling probability for class $c$, $p_{c}$ as:
$p_{c}=\frac{N_{c}}{\sum_{c^{\prime}}N_{c^{\prime}}}.$ (7)
Our Modification. In our method, we initially sample the class category $c$ by using the estimated ID of class $c$, and then uniformly sample from the class samples:
$p_{c}=\frac{\hat{d}_{c}}{\sum_{c^{\prime}}\hat{d}_{c^{\prime}}}.$ (8)
Loss Reweighting methods. Class-based loss reweighting methods scale the loss individually for each class [8, 4]:
$\mathcal{L}=\sum_{c}w_{c}\mathcal{L}_{c},$ (9)
where $\mathcal{L}_{c}$ is the loss for the samples from class $c$ and $w_{c}$ is the loss weight assigned to samples from class $c$. $w_{c}$ is assigned some values based on the cardinality of samples in class $c$; e.g., $w_{c}=\frac{n_{\textrm{min}}}{n_{c}}$ [8], with $n_{\textrm{min}}$ being the cardinality of the least populated class and $n_{c}$ being the cardinality of class $c$.
Our Modification. In our method, we substitute $w_{c}$ by $\hat{d}_{c}\times|\mathcal{C}|$, where $\mathcal{C}$ is the set of classes.
Margin-based methods. Since margin-based methods use different strategies, we integrate $\hat{d}_{c}$ on a case-by-case basis.
In the case of LDAM [9], we use the LDAM + deferred re-weighting framework of [20]. Specifically, LDAM loss has the following structure:
$\mathcal{L}_{\mathrm{LDAM}}(\mathbf{x},y)$ $=-\log\frac{\exp({z_{y}-\Delta_{y}})}{\exp({z_{y}-\Delta_{y}+\sum_{c\neq y}\exp({z_{c}})})},$
$\Delta_{c}$ $=\frac{C}{N_{c}^{\frac{1}{4}}},$ (10)
where $z_{c}$ denotes the output of the model for class $c$, $\Delta_{c}$ is the margin for class $c$, and $C$ is a constant.
Our Modification. Similar to Baltaci et al. [41], we use the ID estimate for each class by setting the margin as:
$\Delta_{c}=0.5\times\frac{\hat{d_{c}}}{\max_{c}\hat{d_{c}}}$. (11)
DRO-LT [10] adjusts the margin with a more robust measure:
$\mathcal{L}_{\textrm{DRO}}(\mathbf{x},y)$ $=-w(y)\log\frac{\exp({-d(\hat{\mu}_{y},\mathbf{x})-\Delta_{y}})}{\sum_{\mathbf{x}^{\prime}}\exp({-d(\hat{\mu}_{y},\mathbf{x}^{\prime})-\Delta^{\prime}_{y}})},$
$\Delta_{y}$ $=2\varepsilon_{y},$ (12)
where $d(\hat{\mu}_{c},z)$ is the distance in the feature space between the sample $\mathbf{x}$ and the estimated centroid of its class $\hat{\mu}_{y}$, and $\Delta_{y}$ is the class margin for class $y$.
Our Modification. Similar to the case of LDAM loss, we utilize class ID estimates by assigning the margin as:
$\Delta_{c}=\frac{\hat{d_{c}}}{\sum_{c^{\prime}}\hat{d_{c^{\prime}}}}\times C.\\$ (13)
In the case of learnable margins in DRO-LT Loss, we initialize the margin values $\varepsilon_{c}$ with the normalized ID estimates $\hat{d}_{c}$ and allow these epsilon values to update during training as done in the original paper.
The Post-hoc logit adjustment method proposed by Menon et al. [3], predicts a label according to:
$\operatorname*{argmax}_{y}\frac{\exp(w_{y}^{\top}\Phi(x))}{\pi_{y}^{\tau}}=\operatorname*{argmax}_{y}f_{y}(x)-\tau\log\pi_{y},$ (14)
where $f_{y}(x)=w_{y}^{\top}\Phi(x)$ represents the logits of a neural network, $\tau$ a tuning parameter and $\pi\in\Delta_{y}$ are estimates of the class priors $p(y)$.
Our Modification. Instead of using the empirical class frequencies for $\Delta_{y}$, we adopt $\hat{d}_{y}$ as follows:
$\Delta_{y}=\frac{1/\hat{d}_{y}}{\sum_{c^{\prime}}1/\hat{d}_{c^{\prime}}}.$ (15)
6 Experiments and Results
Datasets. We evaluate our method through the CIFAR-LT(10/100), Places-LT and ImageNet-LT datasets that are commonly used in the long-tailed visual recognition literature, as well as the semantically imbalanced SVCI-20 dataset. For details about the datasets, see C.1.
Evaluation Measure. As in prior work, we use the top-1 classification accuracy: $N^{v}_{t}/N^{v}$, where $N^{v}$ is the number of examples in the validation set, and $N^{v}_{t}$ is the number of correctly-predicted examples.
Training & Implementation Details are provided in C.2.
Dataset: CIFAR-10-LT CIFAR-100-LT
Imbalance Ratio ($\rho$): 100 50 100 50
Reweighting-based methods
CE (baseline) 71.6 76.3 38.9 42.9
CB + Focal 73.9 79.5 38.4 43.1
CB + Focal [2S] 74.4 78.8 38.0 44.5
CSCE [8] [2S] 74.7 79.2 41.1 45.4
CB + Focal + ID [Ours] 74.8 78.7 40.8 43.0
ID [Ours] [2S] 76.4 80.8 41.7 46.2
Resampling-based methods
CE (baseline) 71.6 76.3 38.9 42.9
PB [6] 72.7 77.8 39.1 43.6
PB [6] [2S] 66.5 75.4 38.7 43.1
CB sampling [7] 72.3 77.0 33.4 40.4
CB sampling [7] [2S] 71.0 78.7 40.2 44.3
ID [Ours] 71.1 78.5 35.2 39.4
PB-ID [Ours] 74.4 78.9 39.7 45.1
PB-ID [Ours] [2S] 74.8 78.8 40.8 44.8
BoT [20] [2S] 80.0 83.6 47.8 51.7
BoT + ID [Ours] [2S] 81.8 85.1 48.6 53.1
Table 1: Exp. 1: Top-1 accuracy results for plug-and-play reweighting and resampling methods for ResNet-32 on long-tailed CIFAR datasets. [2S]: Two-stage methods.
6.1 Experiment 1: Model-Free Approaches
We first evaluate the performance of ID-based imbalance mitigation against alternative plug-and-play model-free class imbalance mitigation methods on several datasets.
Re-sampling and Re-weighting. The results on the CIFAR-LT datasets in Table 1 show that both ID-based resampling and ID-based reweighting methods outperform alternative methods. Furthermore, we observe that our method keeps its effectiveness under increasing data imbalance, in our detailed analysis of one-stage methods in D.
Large-scale Datasets. Similarly, our experiments with the Places-LT and ImageNet-LT datasets indicate that ID-based mitigation is also effective with larger-scale datasets (see Table 2). In particular, ID-based resampling scores at least $0.8\%$ higher than the top re-sampling methods on both datasets, while increasing accuracy by at least $1.1\%$ compared to the top re-weighting methods.
Margin Adjustment. Aside from reweighting and resampling, we also experiment with integrating ID into margin-based methods on a case-by-case basis in Table 3 (additional results on CIFAR-100 can be found in D). In this case, ID-based mitigation performs better in some cases while providing second-best results in others.
Summary. Overall, our results indicate that using data ID ($\hat{d}_{c}$) provides better results than cardinality-based approaches and performs on par with learnable margin-based approaches. Being model-free, our measure stands out as an easy-to-use alternative measure of class imbalance.
Dataset: Places-LT ImageNet-LT
Backbone: ResNet-152 ResNet-10
Reweighting-based methods
CE (baseline) 32.68 34.01
Focal Loss [42] 31.10 32.64
CB [4] 39.44 33.50
CB [4] [2S] 40.44 39.27
CSCE [8] [2S] 40.70 39.43
CB + Focal 37.36 31.04
CB + Focal [2S] 41.02 35.88
ID [Ours] 35.35 37.11
ID [Ours] [2S] 40.91 39.87
CB + Focal + ID [Ours] 39.13 37.26
CB + Focal + ID [Ours] [2S] 39.09 38.82
Resampling-based methods
CE (baseline) 32.68 34.01
PB [6] 40.45 38.48
PB [6] [2S] 41.87 39.74
CB [7] 40.01 34.16
CB [7] [2S] 40.79 39.27
ID [Ours] 38.12 33.27
ID [Ours] [2S] 41.65 39.55
PB-ID [Ours] 38.52 38.78
PB-ID [Ours] [2S] 42.62 40.55
BoT [20] 43.07 43.16
BoT + ID [Ours] 43.39 42.92
Table 2: Exp. 1: Top-1 accuracy results for plug-and-play reweighting and resampling methods on Places-LT and ImageNet-LT. Two-stage methods marked with [2S].
Dataset: CIFAR-10-LT
Imbalance Ratio ($\rho$): 10 50 100
CE (baseline) 87.00 76.29 71.58
LDAM + DRW [9] 87.76 81.64 77.19
Log. Adj. [3] - - 77.30
DRO-LT (learned $\epsilon$) [10] 91.02 85.88 82.39
LDAM + DRW + ID [Ours] 87.81 81.12 76.39
Log. Adj. + ID ($\tau=\tau*$) [Ours] - - 74.10
DRO-LT + ID ($\varepsilon=$ N. ID) [Ours] 90.56 85.39 81.73
DRO-LT + ID ($\varepsilon=$ L. + ID) [Ours] 91.10 85.91 82.34
Table 3: Exp. 1: Top-1 accuracy results for simple plug-and-play margin-based methods on long-tailed CIFAR-10. N.: normalized. L: Learned epsilon. Log. Adj.: Logit Adjustment. See D for the results on CIFAR-100-LT.
6.2 Experiment 2: SOTA Comparison
In these experiments, we integrate ID into state-of-the-art (SOTA) methods and evaluate its performance. Following the success of ID-based resampling in Experiment 1, we integrate ID-based resampling to the methods in these experiments with virtually no additional cost.
The results in Table 4 on CIFAR-LT suggest that, with Bag of Tricks [20] + ID, we boost the performance of the original model in all categories, while our GLMC [43] + ID method surpasses the state-of-the-art for CIFAR-LT classification in two of the four long-tail categories. In the case of Places-LT (Table 5), ID-based resampling combined with Bag of Tricks [20] again achieves the best accuracy with a low variance. On ImageNet-LT, we compare methods using ResNet-50 and ResNeXt-50 backbones following prior work and show the results in Table 7. We see that ID-based resampling boosts the performance of both Bag of Tricks [20] and BCL [44] with ResNet-50 backbones, with the latter achieving state-of-the-art results for a single model under 100 epochs. In the case of ResNeXt-50 backbones, BCL [44] + ID again improves the accuracy of the original BCL method, and obtains the best accuracy under 100 epochs.
Summary. Our results indicate that data ID can complement and improve existing imbalance mitigation strategies, achieving SOTA results on several datasets.
Dataset: CIFAR-10-LT CIFAR-100-LT
Imbalance Ratio ($\rho$): 100 50 100 50
CB + Focal - 74.6 78.8 38.0 44.5
BBN [45] CVPR’20 79.8 82.2 44.1 49.2
BALMS [23] NeurIPS’20 84.9 - 50.8 54.2
BoT [20] AAAI’21 80.0 83.6 47.8 51.6
MiSLAS [46] CVPR’21 82.1 85.7 47.0 52.3
BCL [44] CVPR’22 84.3 87.2 51.9 56.6
GCL [47] CVPR’22 82.7 85.5 48.7 53.6
TSC [48] CVPR’22 79.7 82.9 43.8 47.4
ResLT [25] TPAMI’22 82.4 85.2 48.2 52.7
GLMC [43] CVPR’23 87.8 90.3 57.5 62.5
MDCS [49] ICCV’23 85.8 89.4 53.2 57.2
GPaCo [50] TPAMI’23 - - 52.3 56.4
SURE + RW [51] CVPR’24 86.9 90.2 57.3 63.1
BoT + ID [Ours] - 81.8 85.1 48.6 53.1
SURE + RW + ID [Ours] - 87.0 90.4 57.7 62.7
GLMC + ID [Ours] - 86.8 90.6 57.6 62.8
Table 4: Exp. 2: Top-1 accuracy results for state-of-the-art methods on long-tailed CIFAR datasets with a ResNet-32 backbone. denotes results reproduced using the code released by the authors.
Method Acc.
OLTR [52] CVPR’19 35.9
BALMS [23] NeurIPS’20 38.7
BoT [20] AAAI’21 43.1
DisAlign [22] CVPR’21 39.5
SADE [27] (3 experts) NeurIPS’22 40.9
ResLT [25] TPAMI’22 39.8
MDCS [49] ICCV’23 42.4
GPaCo [50] TPAMI’23 41.7
APA + AGLU [26] ECCV’24 42.0
PB-ID [Ours] - 42.6
BoT + ID [Ours] - 43.4
Table 5: Exp. 2: Top-1 accuracy results for state-of-the-art methods on Places-LT with a ResNet-152 backbone.
6.3 Experiment 3: ID and Semantic Imbalance
We now consider a scenario with “semantic imbalance”, where classes have the same cardinalities but different semantic difficulties. With cardinalities being the same across classes, this form of imbalance is impossible to address with cardinality-based imbalance mitigation methods. To this end, we use a semantically-imbalanced natural image dataset, SVCI-20 [41], containing data from 20 classes – 10 classes each from the CIFAR-10 and SVHN datasets, respectively. The dataset is processed such that all classes have the same number of training samples (by random sampling $N$ examples per class), leaving semantic imbalance as the prominent form of imbalance in the dataset.
As shown in Figure 5, the individual ID estimates of CIFAR-10 classes are much higher than SVHN classes, indicating a clear separation of CIFAR-10 classes from SVHN classes within the SVCI-20 dataset when using ID-based model training. This shows that our ID-based approach does capture imbalance among classes where cardinality-based methods fail. Individual ID values for SVCI-20 classes can additionally be found in D.4.
Imbalance mitigation results in Table 6 indicate that our ID-based approach provides better results and is not affected by cardinality, unlike the cardinality-based approaches.
Summary. We demonstrate that cardinality-based methods fail in datasets with classes of equal cardinality but different semantic complexity. However, our ID-based approach is effective in such semantic imbalance scenarios.
$N=250$ $N=100$
CE / Cardinality-based methods [*] 77.96 $\pm$ 0.65 66.44 $\pm$ 1.40
Focal Loss [42] 78.35 $\pm$ 0.04 66.39 $\pm$ 0.97
PB-ID [Ours] 79.02 $\pm$ 0.18 67.29 $\pm$ 0.84
Table 6: Exp. 3: Top-1 balanced accuracy values for plug-and-play resampling and reweighting methods on the balanced SVCI-20 dataset. NN: Number of samples per class. [*]: [4, 8, 7, 6].
Figure 5: Exp. 3: Histogram of estimated ID values of the two semantically-different types of classes in the SVCI-20 dataset. See D.4 for all ID values.
6.4 Experiment 4: Ablation Analysis
We now evaluate our ID-based approach with different ID estimators: FisherS [29], maximum likelihood estimator (MLE) [32] with asymptotic correction [53], and Tight Locality Estimation (TLE) [33]. For comparison on CIFAR-LT, we use the GLMC [43] and SURE + re-weighting methods [51] (Table 8), and for ImageNet-LT, we use BCL [44] – see Appendix for all results.
Summary. Overall, we see that our approach works with other ID estimators and that there is no single ID estimator which is best for all methods. With SURE + ID and BCL + ID, our method equipped with the FisherS estimator works best, while with GLMC + ID we observe that our method with TLE and MLE provides better results.
Category: Many Med. Few Total
Backbone: ResNet-50, Training: $\leq 100$ epochs
LDAM [9] NeurIPS’19 60.4 46.9 30.7 49.8
BoT [20] AAAI’21 - - - 50.8
Logit Adj. [3] ICLR’21 61.1 47.5 27.6 50.1
RIDE [54] (3E) ICLR’21 - - - 54.9
MiSLAS [46] CVPR’21 62.9 50.7 34.3 52.7
ACE [55] (3E) ICCV’21 - - - 54.7
BCL [44] CVPR’22 - - - 56.0
GCL [56] CVPR’22 63.0 52.7 37.1 54.5
TSC [48] CVPR’22 63.5 49.7 30.4 52.4
BoT + ID [Ours] - - - - 51.3
BCL + ID [Ours] (90EP) - 66.6 53.8 37.6 56.5
Backbone: ResNeXt-50, Training: $\leq 100$ epochs
BALMS [23] NeurIPS’20 52.2 48.8 29.8 51.4
RIDE [54] (3E) ICLR’21 68.0 52.9 35.1 56.3
DisAlign [22] CVPR’21 62.7 52.1 31.4 53.4
ACE [55] (3E) ICCV’21 71.7 54.6 23.5 56.6
BCL [44] CVPR’22 67.6 54.6 37.5 57.3
ResLT [25] TPAMI’22 63.6 55.7 38.9 56.1
BCL + ID [Ours] - 69.0 54.7 37.8 57.9
Backbone: ResNeXt-50, Training: $\leq 180$ epochs
BALMS [23] NeurIPS’20 65.8 53.2 34.1 55.4
PaCo [57] (180EP) ICCV’21 64.4 55.7 33.7 56.0
BCL [44] CVPR’22 - - - 57.1
SADE [27] (3E) NeurIPS’22 67.0 56.7 42.6 58.8
GLMC [43] (135EP) CVPR’23 70.1 52.4 30.4 56.3
DSB + RIDE [17] ICLR’23 68.6 54.5 38.5 58.2
MDCS [49] (3E) ICCV’23 - - - 60.2
GLMC + ID [Ours] - 66.8 53.7 34.4 56.3
BCL + ID [Ours] - 70.9 53.9 37.2 58.2
Table 7: Exp. 2: Top-1 accuracy results for state-of-the-art methods on ImageNet-LT under 100 and 180 epochs. denotes methods that use RandAugment [58] augmentation. 3E: 3 experts. EP: Epochs.
Dataset: CIFAR-10-LT CIFAR-100-LT
Imbalance Ratio ($\rho$): 100 50 100 50
SURE + RW [51] 86.9 90.2 57.3 63.1
SURE + RW + ID (w/ MLE) 86.2 90.0 54.8 59.7
SURE + RW + ID (w/ TLE) 86.0 90.0 56.0 60.8
SURE + RW + ID (w/ FisherS) 87.0 90.4 57.7 62.7
GLMC [43] 87.8 90.3 57.5 62.5
GLMC + ID (w/ MLE) 87.1 90.7 57.2 63.0
GLMC + ID (w/ TLE) 87.9 90.5 58.0 62.8
GLMC + ID (w/ FisherS) 86.8 90.6 57.6 62.8
Table 8: Exp. 4: Top-1 accuracy results for ID-based methods with different ID estimators on CIFAR-LT. denotes results reproduced using the code released by the authors.
7 Conclusion
Prior work addresses imbalance by relying on class cardinality or by training multi-stage, complex models. In this paper, we proposed “data intrinsic dimensionality (ID)” as an alternative measure for imbalance. We show that Data ID can measure imbalance in a model-free, training-free, and cardinality-independent way, making it highly practical and versatile for a range of imbalance mitigation methods. Our experiments show that Data ID achieves results that are on par with or better than state-of-the-art methods across several datasets with varying degrees of cardinality imbalance.
Acknowledgments
This work was supported by the Council of Higher Education Research Universities Support program as a Middle East Technical University (METU) Scientific Research Projects project (“New Techniques in Visual Recognition”, Project No. ADEP-312-2024-11485). We also gratefully acknowledge the computational resources provided by METU-ROMER, Center for Robotics and Artificial Intelligence.
Appendices
Appendix A Extended Literature
A.1 Literature on Long-Tailed Visual Recognition
Long-tailed recognition approaches can be broadly analyzed under the following main categories [2, 1, 20]:
Resampling Methods modify the sampling probabilities of examples [5], mainly focusing on under-sampling majority classes or over-sampling minority classes. Notably, class-balanced (uniform) sampling [7, 6] gives each class an equal probability of being sampled. To overcome the drawback of overfitting to tail classes in a class-balanced setting, various methods are proposed, such as soft-balancing [59], progressive balancing [6], and context-based augmentation [60].
Loss Reweighting Methods adjust the weights of the objective function during training to prioritize less representative classes. Many approaches propose reweighting strategies relying on class frequency [8, 9, 21, 22]. A prominent example by Cui et al. [4] is the class-balanced loss, which computes loss weights using the effective number of samples for each class. Similarly, Park et al. [61] accounts for the influence [62] of samples on decision boundary for reweighting. Another well-known example, Focal loss [42], applies loss reweighting to each example depending on the relative difficulty (based on its prediction probability) of that example during training. There also exist methods that leverage meta-learning [63, 64, 65] to estimate loss weights from a balanced meta-data.
Margin (Logit) Adjustment Methods enforce margins in logit space to obtain a balanced decision boundary [23, 24]. Common approaches [9, 3, 47] define these margins over class distributions. LGLA [66] applies logit adjustment to experts in an ensemble for balanced global and local margins. Differing from these, DRO-LT [10] enforces a margin in feature space to learn representative features.
Other Approaches with alternative perspectives are also present: E.g., residual fusion of specialized heads [25], exploration of an optimal combination of existing techniques [20, 51], a specialized activation function [26] or distribution-aware expert learning [49, 54, 55, 27]. Moreover, model calibration techniques to overcome over-confidence show promising improvements [46, 22]. Complementarily, a branch of works focuses on learning unbiased feature representations, with supervised contrastive learning [57, 44, 50, 48] or data augmentation [43].
While this paper presents a brief introduction, there are other families of methods not addressed in this section. For a comprehensive overview, we refer readers to [2, 1, 20].
A.2 Literature on Intrinsic Dimension Estimation
Intrinsic dimensionality analysis focuses on quantifying the intrinsic qualities of data. As a general definition, the intrinsic dimension is the minimum number of parameters required to represent a data distribution with minimal loss of generality [11]. ID is closely related to the manifold hypothesis [14], where it is assumed that natural image datasets with high dimensionality lie on a manifold of lower dimensionality, which makes learning feasible. In this context, ID refers to the number of dimensions of this lower-dimensional manifold. The ID of a dataset can be estimated globally (i.e., for the entire dataset), or locally (around a local neighborhood), depending on the scope and the estimation method. ID estimation methods in the literature can be studied in two general categories: Projective methods [28, 29, 30, 31] and geometric (topological) methods [32, 33, 34].
Projective methods project data points into lower-dimensional spaces and minimize the projection error in the resulting subspace to estimate the intrinsic dimension of the data. Linear projection with some variant of Principal Component Analysis (PCA) [67, 36, 28] is a common approach that has been used to estimate ID with varying levels of success. FisherS [29] is another projective method that combines PCA with Fisher separability analysis to determine the intrinsic dimension. Other nonlinear projective methods such as LLE [30] and Isomap [31] have also been used for ID estimation as well.
Geometric (topological) methods focus on geometric manipulation of the data to estimate the ID of the underlying manifold. One popular branch of geometric methods is nearest-neighbor methods, where ID is determined from the nearest neighbors of each point in the dataset [32, 33, 34]. Levina and Bickel [32] propose MLE, which estimates the ID by assigning the sampling probability of point-wise neighbors to a Poisson process and maximizing the associated likelihood, and Facco et al. [34] propose TwoNN, which is another nearest-neighbor method that only requires the two nearest neighbors for each point for ID estimation. Alternatively, fractal-based methods associate the distribution of points in the dataset with a fractal structure, after which tools such as the correlation dimension or the Hausdorff dimension can be used to estimate ID [68].
ID estimators can also combine the methods described above, such as in the work of Ceruti et al. [69], however, there is no clear-cut process to determine the optimal ID estimator. The challenge of selecting an ID estimator comes from the fact that natural image datasets trivially do not have ground truth ID values. For more on ID estimation and evaluation of ID estimators, we refer to [12, 13].
A recent work by Ma et al. [18] focuses on the intrinsic dimension of perceptual manifolds of balanced datasets in the context of model fairness and introduces ID regularization as an end-to-end trainable loss, and similarly the work by Ansuini et al. [35] also studies the ID of the representation of data in the intermediate layers of a neural network. In comparison, we introduce ID as an imbalance measure and show that it can be used for training as a model-free mitigation method in the context of imbalanced datasets.
Appendix B Pseudocode of Our Approach
We provide a high-level algorithm (see Alg. 1) to calculate class-wise IDs for multi-class classification using the FisherS [29] approach, which describes the process underlined in Section 3. Normalized ID estimates computed here can be used with resampling, reweighting and margin-based methods, as described in Section 5.
Algorithm 1 Pseudocode for our ID-based imbalance measure calculation, using FisherS [29]. The estimated class-wise measures are utilized as described by Sec. 5.
1:
[…]
3:$C$: Conditional number.
[…]
6:
[…]
8:for class $c\in\mathcal{C}$ do
9:  # Apply preprocessing steps
10:  Assign $X_{c}\leftarrow\{x\in X\mid\text{$x$ is from class $c$}\}$
11:  Normalize $X_{c}\leftarrow X_{c}-\bar{X_{c}}$
12:  Compute $V,U,S=\mathrm{PCA}(X_{c})$, where $U$ are the projections of $X_{c}$ onto principal vectors $V$ with explained variances $S$ in decreasing order
13:  Select $k$ principal components, where $k=\max\{i:S_{1}/S_{i}<C\}$
14:  Whiten columns of $U$: $u_{i}\leftarrow u_{i}/\sigma(u_{i}),i=1...k$
[…]
16:  # Determine optimal $\alpha$ for class $c$
17:  for $\alpha\in\boldsymbol{\alpha}_{s}$ do
18:   Compute empirical inseparability probabilities $p_{\alpha}$ from the projected data $\mathbf{u}_{1:k}$
19:   Compute empirical mean $\bar{p_{\alpha}}\leftarrow\frac{1}{N}\sum_{i=1}^{N}p_{\alpha}$
20:   Estimate $n_{\alpha}$ from $\bar{p_{\alpha}}$ using Eq. 3
21:  end for
22:  Remove $\alpha$ values from $\boldsymbol{\alpha}_{s}$ with invalid $n_{\alpha}$ estimates
23:  Select $\alpha^{*}\in\boldsymbol{\alpha}_{s}$ closest to $0.9\times\max{\{\boldsymbol{\alpha}_{s}\}}$
24:  Assign ID value $d_{c}\leftarrow n_{\alpha^{*}}$
25:end for
26:Normalize ID values $\hat{d_{c}}\leftarrow d_{c}/\sum_{c^{\prime}}d_{c^{\prime}}$
Appendix C Implementation, Training and Experiment Details
C.1 Dataset Details
CIFAR-LT datasets are long-tailed versions of the standard CIFAR-10 and CIFAR-100 datasets. We adopt the method of Cao et al. [9] in our experiments, following prior work. These datasets follow an exponential long-tailed distribution curve and can be modified by changing the imbalance ratio (IR) of the dataset, with IR defined as:
[…]
In the literature, IR values of 10, 50, and 100 are common for the long-tailed variants of CIFAR-LT datasets.
Places-LT [52] is a long-tailed version of the Places-2 [70] scene recognition dataset. It contains 365 scene classes, with class cardinalities ranging from 5 to 4980 samples per class (with a striking IR of 996). The test set is balanced with 100 samples per class.
ImageNet-LT [52] is a long-tailed modification of ImageNet (ILSVRC-2012) [71]. We use the long-tailed split defined by Liu et al. with 1000 class categories and class cardinalities ranging from 5 to 1280 in the training set. The test set is balanced with 50 samples for each class category.
C.2 Implementation and Training Details
We use ID-based mitigation as a complementary method to existing methods. For this, we use the same pipeline and parameters used in the original methods, only tuning the learning rate accordingly with ID-based reweighting and resampling. For ID estimation, we use the methods from the scikit-dimension [72] library. Our experiments with model-free approaches use the Bag of Tricks [20] framework. In further experiments on ablations and state-of-the-art comparisons, we additionally integrate ID into the GLMC [43], SURE [51] and BCL [44] methods. For our experiments with Places-LT, we finetune a ResNet-152 [73] model provided by torchvision following the prior work in the literature [52, 25, 50].
C.3 Experiment Details
For our experiments using the Bag of Tricks [20] framework, we use the training configurations reported in Tab. 9. For experiments using Focal Loss [42], we additionally sample $\gamma\in\{1.0,2.0\}$.
The complete Bag of Tricks multi-stage method with CAM-based sampling of [20] uses mix-up sampling [74] with $\alpha=1.0$, which we also use in our Bag of Tricks + ID experiments. Following [20], our Bag of Tricks + ID setup consists of a first stage of CE training with default sampling and mix-up augmentation, followed by CAM-based example generation using the trained first stage model. The second stage follows including the CAM-based samples with ID-based resampling and mix-up augmentation, and a short fine-tuning stage with ID-based resampling without mix-up sampling. Following Zhang et al. [20], we double the epochs for training stages that use mix-up augmentation to account for slow convergence.
For our experiments with Places-LT, we finetune a ResNet-152 [73] model provided by torchvision following the prior work in the literature [52, 25, 50]. We train the ResNet-152 backbone with learning rate sampled from $\ell_{b}\in\{.01,.02\}$, and train the classifier (the final fully-connected layer) with a learning rate sampled from $\ell_{c}\in\{.01,.1,.2\}$.
For our experiments with ImageNet-LT, we train with a ResNet-10 backbone from scratch in accordance with the Bag of Tricks [20] framework, and switch to ResNet-50 [73] and ResNeXt-50 [75] based backbones for comparing against the state-of-the-art methods.
Tab. 10 shows the training configuration for BCL [44] + ID on ImageNet-LT. BCL is trained with a base learning rate of 0.1 and a cosine annealing learning rate scheduler. Following [44], we use $\lambda=1.0$ and $\mu=0.35$ as the logit compensation and balanced contrastive learning loss weights respectively.
The training configurations of SURE [51] on the CIFAR-LT datasets are listed in Tab. 11. SURE is trained with sharpness-aware minimization [76] together with stochastic weight averaging [77] (SWA) for increased robustness. SWA is initialized at 120 epochs with a SWA-specific learning rate of 0.05. Mix-up augmentation is used with the mixup parameter sampled from $\mathrm{Beta}(\beta,\beta)$.
Tab. 12 shows the training configuration for GLMC [43] on ImageNet-LT and CIFAR-LT datasets. On CIFAR-LT datasets, the model is trained with the SGD optimizer with a momentum of 0.9 with a cosine annealing learning rate scheduler. Mix-up augmentation is used with the mixing parameter sampled from $\mathrm{Beta}(\beta,\beta)$.
Regardless of the specific backbone, all experiments on CIFAR-LT were conducted on a single RTX3090 GPU, Places-LT experiments on two RTX3090 GPUs, and ImageNet-LT experiments were conducted with four RTX3090 GPUs with torch == 1.12 and scikit-dimension == 0.3.4.
Dataset: CIFAR-LT Places-LT ImageNet-LT
Plug-and-play methods + ID parameters
Learning Rate $\ell\in\{.1,.2,.3\}$ $\ell_{\mathrm{b}}\in\{.01,.02\},$ $\ell\in\{.1,.2\}$
$\ell_{\mathrm{c}}\in\{.01,.1,.2\}$
Epochs 200 30 100
Batch Size 128 128 $\{256,512\}$
Optimizer SGD SGD SGD
Momentum 0.9 0.9 0.9
Weight Decay $2\times 10^{-4}$ $5\times 10^{-4}$ $1\times 10^{-4}$
LR scheduling Step Step Step
LR reduction factor .01 .01 .01
LR reduction epochs $\{120,160\}$ $\{16,24\}$ $\{60,80\}$
Data transforms (Train time) Random resized cropping, horizontal flipping, color jitter
Data transforms (Test time) Resizing and center cropping
Bag of Tricks + ID additional parameters
LR (1S) .1 $\ell_{b}=.001,\ell_{c}=.01$ $\ell\in\{.1,.2\}$
Epochs (1S) 320 40 160
LR scheduler (1S) - multi-step multi-step
LR scheduler epochs (1S) - $\{25,35\}$ $\{100,140\}$
LR reduction factor (1S) - .1 .1
LR (2S) $\ell\in\{.0003,.001\}$ $\ell_{b}=.001$, $\ell\in\{.002,.003\}$
$\ell_{c}\in\{.001,.01\}$
Epochs (2S) 80 20 40
LR scheduler (2S) multi-step multi-step multi-step
LR scheduler epochs $\{40\}$ $\{8,16\}$ $\{30\}$
LR reduction factor .01 .1 .1
LR (fine-tuning) $\ell\in\{.0003,.001\}$ $\ell_{b}=\{.0003,.001\}$, .002
$\ell_{c}\in\{.0003,.001\}$
Epochs (fine-tuning) 20 8 20
LR scheduler epochs $\{8,16\}$ $\{4,6\}$ $\{8,16\}$
LR reduction factor .1 .1 .1
Table 9: Training configurations for experiments using the Bag of Tricks [20] framework. indicates training using mixup [74] augmentation with $\alpha=1.0$, where we follow [20] in doubling the number of epochs to account for slow convergence. LR: Learning Rate.
Dataset: ImageNet-LT
Learning Rate 0.1
Batch Size 256
Epochs $\{90,180\}$
Cosine classifier
Weight Decay $5\times 10^{-4}$
Contrastive Views RandAug - RandAug
Contrastive Learning Temperature 0.07
Optimizer SGD
Momentum 0.9
LR Scheduler Cosine annealing
Table 10: Training configuration for BCL [44] + ID on ImageNet-LT.
Dataset: CIFAR-LT
First stage training
Learning rate 0.1
Batch size 128
Epochs 200
Weight Decay $5\times 10^{-4}$
Optimizer SAM (SGD)
Cosine classifier
$\lambda_{\mathrm{crl}}$ 0.0
$\lambda_{\mathrm{mixup}}$ 1.0
$\beta$ 10
Second stage fine-tuning
Learning rate 0.005
Batch size 128
Epochs 50
Re-weighting type exponential
Cosine classifier
$\lambda_{\mathrm{crl}}$ 0.0
$\lambda_{\mathrm{mixup}}$ 1.0
$\beta$ 10
Table 11: Training configurations for SURE + RW [51] + ID on CIFAR-LT.
Dataset: ImageNet-LT CIFAR-10-LT CIFAR-100-LT
($\rho=100/50$) ($\rho=100/50$)
Learning rate 0.1 0.01 0.01
Batch Size 120 64 64
Epochs 135 200 200
Beta 0.5 0.5 0.5
Momentum 0.9 0.9 0.9
Weight Decay $2\times 10^{-4}$ $5\times 10^{-3}$ $5\times 10^{-3}$
Contrast Weight 10 1 4 / 6
Label Weighting 1.0 1.2 1.2
Table 12: Training configuration for GLMC [43] + ID on CIFAR-LT and ImageNet-LT.
Appendix D Additional Results
Dataset: CIFAR-10-LT CIFAR-100-LT
Imbalance Ratio ($\rho$): 10 50 100 10 50 100
CE (baseline) 87.00 76.29 71.58 56.04 42.89 38.90
LDAM + DRW [9] 87.76 81.64 77.19 57.57 46.44 41.83
Log. Adj. [3] - - 77.30 - - 44.10
DRO-LT (learned $\epsilon$) [10] 91.02 85.88 82.39 64.02 53.95 48.57
LDAM + DRW + ID [Ours] 87.81 81.12 76.39 57.23 45.38 41.46
Log. Adj. + ID ($\tau=\tau*$) [Ours] - - 74.10 - - 41.00
DRO-LT + ID ($\varepsilon=$ N. ID) [Ours] 90.56 85.39 81.73 63.40 53.12 48.52
DRO-LT + ID ($\varepsilon=$ L. + ID) [Ours] 91.10 85.91 82.34 63.69 53.63 48.52
Table 13: Experiment 1: Top-1 accuracy results for simple plug-and-play margin-based methods on long-tailed CIFAR datasets. N: normalized. L: Learned epsilon. Log. Adj.: Logit Adjustment.
D.1 Rarity-based Analysis on One-Stage Model-Free Approaches
We further analyze the results of ID-based imbalance mitigation with one-stage model-free methods in a rarity-based grouping and present the results in Tab. 14. We sort the 100 classes of CIFAR-100 by decreasing example counts and group the 33/34/33 class partitions as head, middle, and tail classes respectively. ID-based reweighting and resampling both outperform alternative methods, and in particular the results show that ID-based mitigation primarily increases the performance of trained in the middle and tail classes of a dataset.
Imbalance Ratio ($\rho$): 50 100
Rarity: Head Middle Tail Average Head Middle Tail Average
Reweighting-based methods
CE (baseline) 67.7 42.7 18.3 42.9 65.9 39.1 11.7 38.9
Focal Loss [42] 68.4 42.5 15.1 42.0 66.9 37.4 9.4 37.9
CB + Focal 65.2 44.1 20.0 43.1 67.6 37.4 10.3 38.4
ID [Ours] 65.8 43.6 19.3 42.9 66.3 38.2 12.1 38.9
ID + CB + Focal [Ours] 65.3 44.5 19.0 43.0 66.8 41.6 14.1 40.8
Resampling-based methods
CE (baseline) 67.7 42.7 18.3 42.9 65.9 39.1 11.7 38.9
CB sampling [7] 61.6 41.3 18.1 40.4 57.2 32.5 10.4 33.4
PB sampling [6] 66.2 45.1 19.3 43.6 64.9 40.3 12.1 39.1
ID [Ours] 52.5 39.7 18.3 36.9 50.8 33.4 10.2 31.5
PB-ID [Ours] 65.8 47.0 22.3 45.1 64.9 40.2 14.1 39.7
Table 14: Experiment 1: Top-1 accuracy results for one-stage plug-and-play resampling and reweighting methods for ResNet-32 on long-tailed CIFAR-100.
D.2 Additional Analysis on FisherS ID Estimation
FisherS [29] calculates data ID by estimating the mean inseparability of the normalized dataset and determining the candidate dimension for the data which best matches this degree of inseparability for given parameter $\alpha$. On the equidistribution to a sphere, Fig. 6(a) shows that, as the dimension of the sphere increases, the mean inseparability $\overline{p_{\alpha}}$ on the sphere decreases rapidly following the blessing of dimensionality [14]. Similarly, given a parameter $\alpha$ and estimated mean inseparability $\overline{p_{\alpha}}$, the estimated intrinsic dimension $n_{\alpha}$ again decreases sharply with increasing inseparability as shown in Fig. 6(b).
(a) Average inseparability $\overline{p_{\alpha}}$ against increasing dimension $n$.
(b) Estimated $n_{\alpha}$ against average inseparability $p_{\alpha}$.
Figure 6: Interactions between estimated dimension and degree of separability in FisherS.
In addition to the analysis of FisherS as an imbalance measure (in Section 3), in Fig. 7 we plot the ID estimates of classes in the CIFAR-10-LT dataset under increasing cardinality imbalance. We again observe that data ID estimates remain relatively robust under the change in the imbalance ratio.
Figure 7: ID estimations of FisherS [29] under increasing class imbalance for CIFAR-10-LT.
Finally, in Fig. 8 we analyze the sensitivity of our method equipped with the FisherS estimator to the choice of $C$, controlling the number of components selected in dimensionality reduction. Our findings indicate that our method is robust to changes to $C$ in synthetically generated datasets and the CIFAR-10 dataset, respectively.
(a) ID estimates with different $C$ values on synthetic multivariate Gaussian data.
(b) Normalized ID estimates for CIFAR-10 with uniformly sampled $C$ values from $[4,16]$.
Figure 8: The sensitivity of ID estimation to $C$.
Figure 9: FisherS ID estimates with different hyperparameters on CIFAR-10.
D.3 Additional Ablation Results on the Choice of the ID Estimator
Along with the results reported in Section 6.4, we additionally perform an ablation for the choice of ID estimator with the ImageNet-LT dataset in Tab. 15, where the results indicate that the FisherS estimator performs better than alternative estimators when paired up with the BCL [44] method.
In general, our results indicate that there is no singular optimal choice of ID estimator for all cases, and the optimal estimator may be different for a particular dataset and a particular method.
Dataset: ImageNet-LT
BCL [44] 56.1
BCL + ID (w/ MLE) 54.4
BCL + ID (w/ TLE) 55.5
BCL + ID (w/ FisherS) 56.5
Table 15: Experiment 4: Top-1 accuracy results for ID-based methods with different ID estimators on ImageNet-LT with a ResNet-50 backbone trained for 90 epochs.
D.4 Additional Robustness Analysis of Our Method
Analysis wrt. Dimension Correlation. We investigate the effect of increasing variance and correlation between the intrinsic dimensions of the data. For ID = $d$, we sample from multivariate normal distributions $\mathcal{N}(\boldsymbol{\mu},\mathbf{\Sigma})$ where $\boldsymbol{\mu}$ is a $d$-dimensional zero vector and $\mathbf{\Sigma}$ is a $d\times d$ covariance matrix. In our analysis of the effect of changing variance, we tailor our methods according to the type of covariance matrix:
  • 1.
    For normal distributions with spherical covariance matrices ($\mathbf{\Sigma}=\sigma I$), we change the scaling factor $\sigma$ and observe that ID estimation is robust against changing variance, as seen in Fig. 10(a).
  • 2.
    For diagonal covariance matrices ($\mathbf{\Sigma}=\mathrm{Diag}(\boldsymbol{\sigma}_{1,\dots,N})$), we generate diagonal covariance matrices with fixed total variation ($\mathrm{var}_{t}=\mathrm{trace}(\mathbf{\Sigma})$) and observe the change in ID with increasing total variation. As Fig. 10(b) shows, ID is relatively robust to changing variance in diagonal covariance matrices.
  • 3.
    For full covariance matrices, we generate arbitrary full covariance matrices with fixed generalized variance ($\mathrm{var}_{g}=|\mathbf{\Sigma}|$), and observe ID estimates as we gradually increase the generalized variance. In this case, Fig. 10(c) shows that ID underestimates the actual ID of the data, however, this underestimation is not correlated with increasing generalized variance.
(a) ID estimates with a spherical covariance matrix.
(b) ID estimates with a diagonal covariance matrix.
(c) ID estimates with a full covariance matrix.
Figure 10: ID estimation on synthetic data with known dimensionality with changing variance. Dashed lines indicate the ground truth ID values.
D.5 Failure Cases with Our Method
We also briefly explore cases where our ID-based method may fail to be effective in imbalance mitigation. Some examples of failure cases for our approach is when ID fails to be informative for a particular class; one way to assure this is to reverse or shuffle the ID values for each class in the dataset. Tab. 16 shows the result of mitigation with uninformative ID values on CIFAR-10, where the results trail behind standard CE-based training.
Dataset: CIFAR-10 CIFAR-100
CE 71.60 38.90
CSCE [8] 74.69 41.08
ID 76.44 41.65
ID (Reversed) 71.55 39.33
ID (Shuffled) 71.52 $\pm$ 0.85 38.84 $\pm$ 0.61
Table 16: Top-1 accuracy results with uninformative ID configurations for CIFAR datasets with an imbalance ratio of 100.
Another possible failure case is in the extreme case where there are not enough samples to reliably estimate ID for a class; for example, estimating ID for a class with just one sample is not feasible. In particular, Fig. 11 shows that the variance in estimated ID increases with very low sample sizes, which indicates that ID estimates for the tail-most classes in datasets with extreme imbalance might be uninformative due to high variance.
Figure 11: ID estimates with low sample sizes on CIFAR-10.
CIFAR-10 Class ID SVHN Class ID
airplane 8.82 0 4.78
automobile 10.83 1 4.65
bird 6.59 2 4.71
cat 10.65 3 2.98
deer 8.75 4 3.78
dog 11.56 5 2.92
frog 9.06 6 2.86
horse 11.74 7 3.92
ship 9.78 8 5.39
truck 13.57 9 4.67
Table 17: Experiment 3: ID estimates for classes in the balanced SVCI-20 dataset with $N=100$ samples per class.